Use the substitution method to solve the system of equations. Enter your answer as an ordered pair.
y = 5x - 17
y = x + 3

Answer by mananth(16946)

-5 x + y = -17 .............1
1 y = 3 + x ---------2
y = 3 + x
Substitute y in (1)
-5x+(3+x)=-17
-5x+3+x =-17
-5x+x=-17-3
-4x=-20
/-4
x= 5
Plug the value of x in (1)
-5 x + y = -17
-5 * 5 + y = -17
-25 + y = -17
y = 8
/ 1
y= 8
(5,8)